OpenStudy (anonymous):

first five terms of an=an-1+6

OpenStudy (anonymous):

Your series is a recursive one. This means that to know the nth term you need to know the vale of the (n-1)th term. Thus to know the second term you need to know the first term. Is the value of the first term given in the question?

OpenStudy (amistre64):

as an aside to terminology. a sequence is a list of elements, either random or generated: listing terms define a sequence. a1, a2, a3, a4, a5 a series is the sum of the terms in a list ... a1 + a2 + a3 + a4 + a5 a sequence of partial sums kinda blends the 2 concepts s1 = a1 s2 = a1 + a2 s3 = a1 + a2 + a3 etc ...

OpenStudy (amistre64):

but as is, you can still list the terms, you just have to leave a1 variable